find the hypotenuse of a right triangle with sides 4in and 3in .
how do yu do this ?

Answer by mananth(16946)

side 1 = 4 in
side 2 = 3 in
using pythagoras theorem
side1^2 +side2^2= hypotenuse ^2
3^2 +4^2 = hypotenuse^2
9+16= hypotenuse^2
25 = hypotenuse ^2
Take square of both sides
5 inch = hypotenuse